### v4

# From Friendship Networks to Classroom Dynamics: Leveraging Neural Networks, Instrumental Variable and Genetic Algorithms for Optimal Educational Outcomes

Abstract
Peer effects significantly shape behavior, attitudes, and performance, particularly in elementary and middle school settings where younger students are highly susceptible to peer influences. By enhancing peer effects, this study uses data from the China Educational Panel Survey (CEPS) to design a classroom assignment policy that maximizes educational outcomes. Our approach comprises three steps: firstly, we develop a friendship formation discrete choice model and estimate it with an interpretable neural network architecture, PeerNN, generating an adjacency-probability matrix $\Omega$ that reflects friendship formation probabilities. In our empirical study, $\Omega$ aligns with well-known social network characteristics such as gender homophily, the presence of central nodes, and varying popularity levels across student subgroups. We also compare the out-of-sample prediction performance of PeerNN and the linear-in-means model. The former outperforms by a huge margin. Secondly, we incorporate $\Omega$ into a linear-in-means model to estimate peer effects. The peer effect parameter, $\beta$, has a different interpretation from the conventional linear-in-means model and opens up a strategic scope of mean-maximizing classroom assignment policy. By exploiting the conditional random classroom assignment in many Chinese middle schools, we construct a valid instrument to address the endogeneity issue induced by $\Omega$ and consistently estimate $\beta$. Lastly, utilizing the estimates of $\Omega$ and $\beta$, we employ a genetic algorithm (GA) to search for the mean-maximizing class assignment policy. Though the result is much more efficient (i.e. more positive average peer effect) than random classroom assignment (i.e. the current practice in most Chinese middle schools), GA policy is highly inequitable: a small number of students are predicted to experience severely negative peer effects. To balance students’ academic performance with educational equity, we propose a fairness metric and penalize classroom assignment that generates large variances in peer effects. The modified method is called algorithmically fair genetic algorithm (AFGA). AFGA policy is less efficient but much more equitable. We allow user-defined parameters for AFGA such that the school principals can adjust the trade-off between efficiency and equity according to their preferences. This comprehensive three-step framework not only improves class assignments but also offers a solution extendable to other group assignment challenges where peer effects are a consideration.
1 Introduction
Peer effects denote the impact individuals within the same social circles exert on each other, influencing various aspects of behavior, attitudes, and performance. Numerous studies (Feld and Zölitz, 2017; Carrell et al., 2018; Shure, 2021; Huang et al., 2021) have explored peer effects, particularly emphasizing their importance in classroom settings. Within the literature on classroom peer effects, one of the popular regression specifications is the linear-in-means model which assumes homogeneity in spillover effects among cohorts. Though the linear-in-means model is easy to implement and has a clear interpretation, it suffers from two major drawbacks. Firstly, the linear-in-means model assumes a uniform strength of the connection between any pair of individuals. Consequently, it does not shed light on network formation and can hardly capture differential peer effects individuals receive from their cohorts. Though under assumptions specified by Manski (1993), the linear-in-means model successfully identifies some parameters related to the average peer effect, researchers may be interested in the entire distribution of peer effect, or at least some aspects of the distribution that are not fully pinned down by the average. To uncover the distribution of peer effect, consideration of network formation is necessary. Linear-in-means model ignores intricacies of network development and therefore, struggles to provide insights on the distribution of peer effect. Secondly, neglecting the network/friendship development hinders the formulation of effective policies that can address issues arising from peer effects. This deficiency in capturing the nuances of network dynamics poses a challenge when optimizing group or class assignments. Therefore, exploring alternative models that account for the complexities of network formation is imperative for more comprehensive learning about peer effect and more effective design of group assignment policy.
In this paper, we expand the linear-in-means methodology by incorporating friendship formation into our model. The modification changes the interpretation of the peer effect parameter and opens up a strategic scope of classroom assignment optimization. Our methodology breaks a school principal’s classroom assignment problem into three parts: predicting friendship formation, measuring peer effect, and formulating an optimal class assignment policy informed by the outcomes of the first two components. Next, we elaborate further on the three-step empirical framework.
First, we predict counterfactual friendship formations under different classroom assignment policies using a content-based approach rather than the conventional graph-based methods. In other words, we exclusively use students’ individual information as predictors, but not their network information as predictors. This approach is necessitated by the absence of linkage information under varying assignments, setting our work apart from mainstream link prediction studies in machine learning (Ahmed and Chen, 2016; Wu et al., 2016; Wang et al., 2014; Dhelim et al., 2022; Chunaev, 2020; Su et al., 2022; Samanta et al., 2021; Ling et al., 2023). Though both our work and these cited papers use neural networks to approximate complicated social network formation processes, we intentionally avoid techniques such as aggregating neighbors’ information which is commonly used in the cited papers. In contrast, we model friendship preferences using a micro-founded discrete choice setup and estimate it with PeerNN, a novel and interpretable neural network design. PeerNN captures social network characteristics such as gender homophily, outperforming traditional linear-in-means models used in peer effect literature (Guryan et al., 2009; Carrell and Hoekstra, 2010; Carrell et al., 2013; Burke and Sass, 2013). Another advantage of PeerNN is that it does not need expensive linkage data as a response variable. Instead, we can use easy-to-collect aggregate relational data (McCormick and Zheng, 2015; Breza et al., 2020; McCormick, 2020).
Second, using the friendship prediction results from PeerNN, we measure peer effects. We replace the restrictive uniform spillover assumption of the linear-in-means model with a more reasonable friendship-weighted specification. This specification leverages the nonuniform friendship formation predicted by PeerNN, addressing the limitations of the linear-in-means model which assumes uniform spillover effects and ignores friendship dynamics (Bramoullé et al., 2009; Goldsmith-Pinkham and Imbens, 2013). However, network formation is endogenous. We exploit the random classroom assignment conditional on school choice to construct an instrument to conduct causal inference. The linear-in-means model measures the causal relationship between classmates’ average characteristics and a student’s outcome, whereas our specification quantifies the causal impact of friends’ weighted average characteristics on a student’s outcome. The change in interpretation of the causal parameter gives rise to a strategic scope for classroom assignment.
Note that there are myriads of channels peer effect can operate through. We focus our study on one specific channel through which peer effects can manifest, namely, the influence of a student’s friends’ average 6th-grade class quantile on the student’s 8th-grade cognitive ability (similar to Murphy and Weinhardt (2020)). The selection of this particular channel is not a central aspect of our research and researchers can readily adapt our framework to investigate alternative channels of peer effect.
In the third step, we optimize classroom assignments in terms of maximizing peer effects. Unlike existing literature, which assumes straightforward dependence between network and group assignment policy (Bhattacharya, 2009; Graham et al., 2014), our problem allows for more intricate dependence between friendship formation and classroom assignment policy. We employ a heuristic optimization method: genetic algorithm (GA) to navigate this complex problem. GA outputs a classroom assignment policy that on average improves peer effect by 1.9% in our counterfactual simulation. However, the output policy is extremely inequitable. In pursuing the highest average peer effect for the whole classroom, GA gives up on a few disruptive peers by predicting that they form an exclusive clique. Consequently, students in that clique experience severely negative peer effects, but overall the entire classroom has a higher average peer effect. Such educational outcomes are extremely inequitable. We propose a modified fairness-aware fitness function for GA and name it algorithmically fair genetic algorithm (AFGA). The new fitness function penalizes the variance of predicted peer effects within and across classrooms. AFGA outputs slightly less efficient (the improvement is 1.2%) but much more equitable policies. This approach aligns with the work of Kleinberg et al. (2018) and Mitchell et al. (2021), who discuss the integration of algorithmic fairness in optimization methods.
This paper has six contributions. (1) We build a micro-founded discrete choice model for friendship formation, the model can be flexibly parameterized and estimated by the neural network method. Our model shows how flexible machine learning algorithms can be adapted to model complicated social processes while maintaining interpretability. The interpretability offers three advantages: (i) we can obtain a friendship probability-adjacency matrix, $\Omega$, as an intermediate product of PeerNN which is an integral part of peer effect estimation and classroom assignment optimization, (ii) the interpretability of $\Omega$ guides the design of loss function of PeerNN which aligns with the discrete choice model even when the response variable is aggregated relational and true linkages are not observed (iii) we can avoid tuning hyperparameters by examining the interpretability of $\Omega$. (2) We modify the classical linear-in-means model by replacing the uniform spillover assumption with a more reasonable friendship-weighted specification. The modified peer effect estimate is interpreted as how much influence a friend has on a student, rather than how much influence a classmate has on a student. The new interpretation opens up a strategic scope of classroom assignment policy optimization whereas the classical linear-in-means fails to do so. (3) Prior research has attempted to conduct causal inference under networked interference (Ma and Tresp, 2021; Jiang and Sun, 2022; Huang et al., 2023), which is exactly the setup in the classroom peer effect. Recognizing that in our data many classrooms are randomly assigned, we disentangle the network interference (counterfactual friendship prediction) and causal inference (peer effect measurement) problems. By exploiting the random assignment, we construct a valid instrument to isolate the peer effect from endogenous friendship selection. (4) Our three-step framework outputs an implementable classroom assignment policy. To the best of our knowledge, this is the first attempt to optimize classroom assignments with real data and consideration of friendship formation. Theoretical works (Berthelon et al., 2019; Bizzotto and Vigier, 2022; Sarkisian and Yamashita, 2024) do not bring their models to real data, whereas empirical works (Bhattacharya, 2009; Graham et al., 2014; Carrell et al., 2013) implicitly assume straightforward link formation like the linear-in-means model (though the peer effect parameter can be heterogeneous) which allows them to solve the optimization problem as mixed integer programming problems (MIPs). Our friendship model is more sophisticated, making the optimization problem not possible to be recast as an MIP. Therefore, we resort to a heuristic optimization method: genetic algorithm (GA). On top of adopting a novel approach (i.e. GA) to the optimization problem, we incorporate common institutional constraints such as gender ratio and classroom size into GA, tailoring it to the group assignment context. (5) While maximizing the average peer effect seems a natural objective of the classroom assignment problem, the plugged-in GA method outputs a policy that generates extremely unfair educational outcomes according to our counterfactual analysis. Therefore, we propose a fairness measure and penalize unfair educational outcomes to balance efficiency and equity. (6) Combining these three elements: predicting friendship formation, measuring peer effects, and optimizing classroom assignments, constitutes a comprehensive framework that extends beyond classroom setups. This framework applies to various group assignment contexts where peer effects are crucial, such as workplace teams (Anderson et al., 2008; Bandiera et al., 2010; Mas and Moretti, 2009), healthcare interventions (Pianese and Belfiore, 2021; Pagkas-Bather et al., 2020), and community programs (Valente et al., 2003; Yu et al., 2024; Cero et al., 2024).
The rest of the paper is organized as follows: Section 2 relates our work to cutting-edge works in various research fields, Section 3 describes the data and introduces the notations that we use in this paper, Section 4 sets up the models and objective functions in our three-step framework, Section 5 demonstrates and explains the results that we obtain from the empirical study and Section 6 concludes the paper and discusses possible future research directions.
2 Literature review
In this section, we summarize how the problems we aim to tackle and the methods we use in this paper relate to and differ from existing literature.
2.1 Classroom assignment
This paper focuses on a specific channel of peer effect: how much a student’s friends’ average 6th-grade class quantile influences the student’s 8th-grade cognitive ability (similar to Murphy and Weinhardt (2020)). We address a principal’s classroom assignment problem: designing a policy rule to optimize the specified peer effect. The closest studies to our paper are Carrell et al. (2013) and Griffith (2024).
In the former paper, the authors conduct an experiment of class assignment aiming to maximize the positive influence of the high-ability students on the bottom 33% students’ GPA. In their study, the experiment result turns out to be surprising as the designed class assignment policy does not improve the bottom 33% students’ performance. The authors reconcile this counterintuitive result by arguing that the linear-in-means model does not consider dynamic friendship formation which aligns with our critique of the linear-in-means model. The first component of our three-step framework addresses this exact problem. We build a micro-founded discrete choice model to capture the friendship formation pattern which is then incorporated into our peer effect estimation and classroom assignment optimization.
The latter paper proves identification results for friendship formation and peer effect measurement, which this paper does not cover. However, to obtain the theoretical results, Griffith (2024) imposes much more restrictive parametric assumptions for friendship formation than PeerNN. Another advantage of PeerNN is that it does not require linkage data like the latter paper, making our paper more broadly applicable and less ethically concerning. In addition, our policy optimization strategy is more systematic. Griffith (2024) simulate a few counterfactual scenarios to compare a small number of assignment policies, our paper adapts GA to search for the optimal policy and subsequently considers algorithmic fairness which will be discussed in Section 2.7. The biggest difference between our paper and Griffith (2024) is that we aim to optimize group assignment policy and the network is policy-dependent, in contrast, the latter paper considers treatment assignment (same as papers cited in Section 2.6) and the network is fixed.
Other attempts to tackle classroom assignment problems or similar group assignment problems include Bhattacharya (2009), Graham et al. (2014), Berthelon et al. (2019), Bizzotto and Vigier (2022) and Sarkisian and Yamashita (2024), we defer the discussion of the difference between these works and ours to Section 2.8 when we compare their choices of optimization method to ours.
2.2 Link prediction for social network using neural networks
The friendship prediction problem in our work differs from the mainstream link prediction problems in social networks (later referred to as the mainstream problem). Daud et al. (2020) summarize that the mainstream problem focuses on (1) predicting future and missing links (Ahmed and Chen, 2016; Wu et al., 2016), (2) recommendation (Wang et al., 2014; Dhelim et al., 2022), (3) detection (Chunaev, 2020; Su et al., 2022), and (4) influence analysis (Samanta et al., 2021; Ling et al., 2023). In contrast, our work predicts counterfactual friendship formation under different classroom assignment policies. Unlike conventional approaches that utilize graph-based measures, we exclusively rely on content-based measures due to the absence of linkage information under different classroom assignments. For the same reason, common methods for analyzing network data, like aggregating neighbors’ information, are inapplicable to our setup. We use example C.1 to illustrate the key differences between our counterfactual friendship prediction problem and the mainstream problem. All examples are collected in Appendix C.
2.3 Modelling friendship formation with aggregate relational data
Our friendship prediction problem entails a counterfactual scenario, wherein we aim to understand how students would select friends under an alternative classroom assignment, operating on the assumption that friendships exclusively develop within the confines of the classroom. Due to the distinctive nature of our research, we are particularly interested in modeling friendship preference. As such, we build a micro-founded economic model that captures friendship preference heterogeneity. Dong et al. (2012); Zhang et al. (2014) adopt a similar concept when building recommendation systems across heterogeneous networks. We estimate the model with a novel and interpretable neural network design: PeerNN. It takes student characteristics as input and one of its intermediate outputs is an adjacency-probability matrix, $\Omega$. Based on our training results, the $\Omega$ matrix modeled by PeerNN captures commonly known social network characteristics including gender homophily, presence of central nodes, and heterogeneous levels of student popularity across distinct subgroups. In terms of network prediction, PeerNN substantially outperforms the linear-in-means model which is the predominant methodology in peer effect literature. Additionally, PeerNN does not require expensive friendship linkage data, instead, its response variable can be easy-to-collect aggregate relational data (McCormick, 2020; Breza et al., 2020; Laga et al., 2021).
2.4 Homophily and transitivity effect in sociology and economics
McPherson et al. (2001); Kossinets and Watts (2009) establish the importance of the homophily effect in network formation. This finding in sociology inspires a large amount of structural work in economics including McCormick and Zheng (2015); Currarini et al. (2016); Breza et al. (2020); Mele (2022). A related concept termed transitivity (i.e. one is more likely to befriend a friend’s friend) has also sparked many structural works in both sociology and economics including Richters and Peixoto (2011); Block (2015); Vasques Filho and O’Neale (2020). By incorporating these two concepts into the loss function of PeerNN, we estimate a structural model for friendship formation.
2.5 Measuring peer effect with linear-in-means model
Traditional studies in peer effect literature, including works by (Guryan et al., 2009; Carrell and Hoekstra, 2010; Carrell et al., 2013; Burke and Sass, 2013), have largely adopted the linear-in-means model, assuming uniform spillover effects among cohorts (See Appendix E for detail of linear-in-means model). This approach, while straightforward and interpretable (Bramoullé et al., 2009; Goldsmith-Pinkham and Imbens, 2013), overlooks the complexity of social networks and fails to capture the varied peer effects individuals experience (Angrist, 2014). By incorporating $\Omega$ into the standard linear-in-means framework, we relax this key assumption prevalent in previous peer effect studies to quantify a peer effect parameter with heightened policy relevance utilizing an instrumental variable approach. The statistically significant and positive peer effect estimate indicates the importance of maximizing the influence of high-achieving students from elementary school to enhance peer effects in middle school classrooms.
2.6 Causal inference under networked interference using neural network
PeerNN + IV + AFGA GNNCE NetEst SPNet
Prediction target Optimal policy Optimal policy Treatment effect Treatment effect
Observable confounder Instrumental variable DDP Adversarial learning DDP
Unobservable confounder Instrumental variable $\times$ $\times$ DDP + treatment prediction
Policy optimization Genetic algorithm Neural network $\times$ $\times$
Table 1: We propose a framework that combines neural network, instrumental variable, and genetic algorithm (PeerNN + IV + AFGA) to solve for optimal class assignment policy. Ma and Tresp (2021); Jiang and Sun (2022); Huang et al. (2023) focus on estimating causal inference using graph neural network (GNN). Their solutions are named GNN-based causal estimator (GNNCE), NetEst, and SPNet respectively. DDP stands for distribution discrepancy penalty.
We summarize the comparison of the past work on causal inference under networked interference using neural network and our work in Table 1. The most salient difference between our model and other methods (Ma and Tresp, 2021; Jiang and Sun, 2022; Huang et al., 2023) in this literature is that while the cited works deal with nonrandom treatment assignment, in our training dataset, students are randomly assigned to different classrooms conditional on observable school choice. Nevertheless, measuring the causal impact of peer effect is not straightforward as friendship network is not randomly assigned. Leveraging on this natural experiment setup of a conditional random classroom assignment, we construct valid instrumental variables to conduct causal inference analysis (Imbens and Angrist, 1995). See Appendix A.1 for a more detailed comparison between our work and this literature.
2.7 Genetic algorithm and algorithmic fairness
The question of optimal classroom assignment has been put forward but left unanswered in Berthelon et al. (2019). Bizzotto and Vigier (2022) and Sarkisian and Yamashita (2024) discuss this problem in a theoretical setup but do not bring it to empirical data. We adopt a genetic algorithm (GA) to solve this combinatorial optimization problem of class assignment with real data (Kumar et al., 2010; Alam et al., 2020). We further include a literature review on evolutionary algorithm in Appendix A.2. Despite GA’s potential benefits, there exists a risk of principals exploiting GA to prioritize maximizing the average peer effect at the expense of students in the lower quantiles. We contribute to the GA literature by considering algorithmic fairness, a concept extensively discussed by Kleinberg et al. (2018); Mitchell et al. (2021). When evaluating fitness in GA, we penalize the standard deviation of predicted peer effect both within and across classrooms. The modified GA termed algorithmically fair genetic algorithm (AFGA), outputs a more equitable class assignment policy. In addition, we demonstrate a trade-off between efficiency and equity when using GA versus AFGA.
2.8 Optimization methods for treatment/group assignment problems
Given the network formation prediction model (PeerNN) and peer effect estimated parameters, a key challenge in solving for the optimal class assignment policy is the dependency of the network structure on the policy itself. This sets our problem apart from the existing literature on treatment assignment under networked interference, which assumes a fixed network (See Appendix A.1). On top of the three papers cited in Section 2.6, there are numerous recent papers that look at the theoretical properties of methods that optimize treatment assignment given a fixed network (Ananth, 2020; Munro et al., 2021; Kitagawa and Wang, 2023a, b; Viviano, 2024). Another stream of literature that solves group assignment problems as either linear programming problems (LP) or mixed integer programming problems (MIP) assumes that link formation between any two individuals is independent of the presence of other individuals. This literature (Bhattacharya, 2009; Graham et al., 2014; Carrell et al., 2013) assumes a friendship formation process like the linear-in-means model, though Bhattacharya (2009); Graham et al. (2014) do allow heterogeneous peer effect parameters. The straightforward dependence between linkage formation and assignment policy allows the optimization problem to be recast as standard programming problems. On the other hand, our friendship formation model accommodates more sophisticated friendship formation patterns such as clustering or more general dependence between a pair of students’ friendship formation and other classmates’ characteristics. Such flexibility makes it impossible for us to recast our optimization problem as an LP or MIP. Due to the added complexity of our problem, we resort to a heuristic optimization method, namely, genetic algorithm (GA), instead of directly solving for an optimal policy, as in the cited works.
3 Data description and data notation
We use the publicly available China Education Panel Survey (CEPS) data provided by the National Survey Research Center at Renmin University of China (NSRC). CEPS conducted surveys on 7,649 students, along with their family members and teachers, across 179 classes. For classes selected to participate, all students were surveyed. We use $N$ to denote the number of students in a single classroom. If it is for a particular classroom $c$ (from school from $s$), we use $N_{c}$ ($N_{cs}$) to denote the class size.
CEPS collected information from students and their family members regarding the students’ 6th-grade details, for example: class rank (in quantile), academic subject interest, extracurricular hobbies, etc. This aspect of the survey is particularly useful for our project. When assigning students to classrooms upon school entry, the principal only has access to their 6th-grade (primary school) information. Therefore, we exclusively use students’ 6th-grade data and other predetermined characteristics (e.g. gender and locality), collectively denoted as $X$, to predict friendship formation. One particular variable in $X$ that we repeatedly utilize in different steps of our methodology is students’ 6th-grade class quantile, denoted as $z$. If we need to emphasize the 6th-grade quantile of all students from classroom $c$ (or a particular student $i$ from classroom $c$ and school $s$), we use $z_{c}$ (or $z_{ics}$) to denote that information.
Students were asked to provide aggregated relational data (ARD) of their friends. The linkage data, though surveyed in section C20 as depicted in Figure 6 in Appendix B, is redacted from CEPS data. Instead, ARD of the friendship network (See section C21 from Figure 6) is provided. Therefore, we have information on ‘out of the five best friends of a student, how many of them                           ?’ where                            can be filled by any question from section C21. We use ARD, denoted as $A_{f}$, as the response variable for PeerNN to uncover friendship formation patterns. $A_{f}$ has size $(N\times 10)$ and records each student’s responses to 10 distinct questions about their friends’ gender, locality, relationship status, etc. Each question only has three possible answers: ‘none’, ‘one or two’, or ‘most of them’. For example, one student may report none of his friends are female, most of his friends are local, and one or two of his friends are in a relationship, etc. Apart from $A_{f}$, each student’s own information regarding these 10 questions is also available in CEPS. The answer to each question is binary (yes or no). For example, one student is female, local, and in a relationship, etc. We denote students’ own information as $A_{s}$, which has the same dimension as $A_{f}$. We use $A_{fiq}$ to denote ARD information provided by student $i$ about question $q$ and $A_{sq}$ to denote all students’ own characteristics for question $q$.
Upon entering their respective secondary schools, 5860 (from 139 classrooms) out of 7649 students (from 179 classrooms) were randomly assigned to classrooms. This natural experiment setup helps generate exogenous variation in data. Exploiting this data-generating property, we construct instrumental variables for causal inference analysis. To keep our input data consistent for all three steps, we use these 5860 students’ data (1) as the training set for PeerNN, (2) to estimate peer effect, and (3) to design classroom assignments. The rest of the 1789 students who are non-randomly assigned to classrooms are used as test data for PeerNN.
4 Model setup
Problem formulation: The ultimate goal of this work is to devise a class assignment policy that maximizes peer effect. The school principal’s objective function is defined as the following
$\max_{\pi\in\Pi}\frac{1}{\sum_{c\in C}N_{c}}\sum_{c\in C}\beta\mathbf{1_{N_{c}}^{\top}}\Omega_{c}(\pi)z_{c}(\pi)$ (1)
where $\pi$ is a policy from a set of policies, $\Pi$, that satisfy mandatory institutional constraints such as balanced gender ratio across classrooms, $C$ is a set of consecutive integers $\{1,2,\dots,|C|\}$ indicating the indices of classrooms in a school and size of $C$ (i.e. $|C|$) is a known fixed number, $\mathbf{1_{N_{c}}}$ denotes a vector of ones of length $N_{c}$, $\{N_{c}\}_{c\in C}$ is a set of integers numbers indicating the prespecified classroom sizes for each classroom $c$, $\Omega_{c}$ is the adjacency matrix of classroom $c$ and $z_{c}$ is the 6th-grade class quantile of all students in the classroom $c$. Both $\Omega_{c}$ and $z_{c}$ are functions of $\pi$ since $\pi$ determines which students are assigned to classroom $c$. $\beta$ is the peer effect parameter: if a student’s friends’ average 6th-grade class quantile improves by 10%, then the student’s cognitive ability test score improves by $\frac{\beta}{10}$ point, holding everything else fixed. We are particularly interested in the sign of $\beta$.
As shown by Figure 1, we solve the class assignment optimization problem with the following three steps: (i) fitting/predicting friendship formation probability, $\Omega_{c}$, using $X$, (ii) measuring peer effect, $\beta$, with an instrumental variable approach, and (iii) heuristically searching a class assignment policy, $\pi$, that maximizes expression (1) with genetic algorithm.
Figure 1: The three-step framework that takes in predetermined characteristics of students and outputs an optimal class assignment policy.
4.1 Friendship formation neural network
We build an interpretable neural network architecture, termed PeerNN, to predict friendship formation with ARD. PeerNN can be divided into four stages: (1) friendship market, (2) linearized propensity score, (3) normalization, and (4) loss function. Next, we describe the four stages along with their interpretation. For notation convenience, the description pertains to a single classroom setting with class size $N$.
Stage 0: Latent feature encoder
We encode our raw data $X$ to a lower dimensional embedding $\sigma(X)$ with a fully connected hidden layer that contains ten nodes, $\sigma(X)$ is an $(N\times 10)$ matrix. This allows students to form friendship preferences based on latent representations of $X$ (i.e. $\sigma(X)$). Stage 0 serves as a feature extraction step for Stages 1-4 and is estimated as part of PeerNN.
Stage 1: Friendship market
Students’ preference parameters $\delta$ (dimension $(N\times 10)$) are functions of $\sigma(X)$. The ith row of $\delta$ matrix, $\delta_{i}$ is the preference parameters of student $i$. $\delta_{ik}$ indicates how much student $i$ prefers to have a friend with a higher value of latent characteristic $k$. Example C.2 illustrates the concept of preference parameters. We use two fully connected layers to model preference parameters $\delta(X)$:
$\delta(X)={\rm ReLu}({\rm ReLu}(\sigma(X)W^{(1)})W^{(2)})\quad\text{where}\quad\sigma(X)={\rm ReLu}(XW^{(0)}).$
Stage 2: Linearized propensity score
We obtain the outer product of $\delta(X)$ and $\sigma(X)$ and denote the outer product as $\Upsilon$, an $N\times N$ matrix, where $\Upsilon_{ij}=\sum_{k=1}^{K}\delta_{ik}\sigma_{jk}=(\delta\cdot\sigma^{\prime})_{ij}$. $\Upsilon_{ij}$ corresponds to a concept in the structural applied microeconomics literature (McFadden, 1974; Train, 2009): the deterministic part of the utility function for student $i$ picking student $j$ as best friend. Alternatively, we can interpret it as the linearized propensity score in causal inference.
Stage 3: Normalization
We manually set $\Upsilon_{ii}$ as negative infinity since students cannot make friends with themselves. Utility for student $i$ to consider student $j$ as a best friend is subject to a utility shock $\xi_{ij}\overset{iid}{\sim}$ Gumbel distribution: $U_{ij}=\Upsilon_{ij}+\xi_{ij}$. Student $i$ picks his best friend based on who gives him the highest utility, $\operatorname*{\arg\!\max}_{j}U_{ij}$. Knowing the joint distribution of $\{\xi_{ij^{\prime}}\}_{j^{\prime}=1}^{N}$, the probability of student $i$ picking student $j$ as best friend is given by softmax function $P(\operatorname*{\arg\!\max}_{j^{{}^{\prime}}}U_{ij^{{}^{\prime}}}=j)=\frac{\exp{(\Upsilon_{ij})}}{\sum_{j^{\prime}}^{N}\exp{(\Upsilon_{ij^{\prime}})}}:=\Omega_{ij}$. The adjacency-probability matrix $\Omega$ is obtained by applying the softmax function on $\Upsilon$ row-wise. Example C.3 in Appendix C demonstrates how Stage 2 and 3 work.
Stage 4: Loss function
PeerNN’s loss function is a weighted sum of three components: (1) fitted value’s MSE, (2) homophily penalty, and (3) transitivity penalty.
Fitted value’s MSE The dependent/response variable of PeerNN is the ARD that students report about their friends $A_{f}$. Let $B_{i}$ denote the number of friends student $i$ reported. We can randomly draw $B_{i}$ friends for $i$th student based on $\Omega_{i}$, summarize his friends’ information about question $q$ with $A_{sq}$, and predict his response about their friends for question $q$, $A_{fiq}$. Our prediction for $A_{fiq}$ is denoted as $\hat{A}_{fiq}(\Omega_{i},A_{sq},B_{i})$. The MSE component of the loss function is $E[(\hat{A}_{fiq}-A_{fiq})^{2}]$. The expectation is taken over all students and questions.
Two issues arise. Due to CEPS’s survey design, an exact mapping from $(\Omega_{i},A_{sq},B_{i})$ to $A_{fiq}$ requires nonlinear operations, rendering MSE to have no closed form. Moreover, sampling friends without replacement based on the probability vector $\Omega_{i}$ is intractable in backpropagation. See appendix D.1 for more details on these issues. To address the two problems, we make two compromises: (1) approximating nonlinear operations with multiple linear functions, and (2) drawing friends with replacement. We can decompose MSE into ${\rm Bias}^{2}$ and ${\rm Var}$. With the two compromises, we derive closed-form expressions for ${\rm Bias}^{2}$ and ${\rm Var}$, exponentially reducing the computational complexities and making the backpropagation process tractable. The description of the two compromises and the derivations of the closed-form expressions are shown in Appendix D.1.
Homophily penalty It is well established in social network formation literature that people tend to make friends who share similarities with themselves. This is called the homophily effect. To incorporate this social network phenomenon into PeerNN’s loss function, we compute the average friend’s profile and penalize the difference between this profile and the students’ own latent characteristics $H(X)=\lVert\Omega(X)\sigma(X)-\sigma(X)\rVert_{2}^{2}$.
Instead of incentivizing PeerNN to predict friends who are similar in raw data $X$, we penalize the difference between average friends’ profile $\Omega(X)\sigma(X)$ and students’ own latent features $\sigma(X)$. The choice between raw data $X$ and latent feature $\sigma(X)$ is not of importance. However, different columns of $X$ do not have the same distributions, for example, gender is a Bernoulli variable whereas class quantile is uniformly distributed. Penalizing directly using $X$ may require us to add a few more weights into the training parameters. Using latent features introduces more flexibility to the homophily penalty and alleviates the need for more weight parameters.
Transitivity penalty By incorporating transitivity into PeerNN, we incentivize the model to form a clustering pattern in predicting friendship. $\Omega_{ij}$ measures the probability of a pair of students $(i,j)$ forming friendship, $(\Omega\Omega)_{ij}$ measures how likely a third student $k$ is considered as best friend by student $i$ and considers student $j$ as best friend. If students $i$ and $j$ are very likely to be best friends, then they are very likely to have a ‘mutual’ best friend in a third student $k$. While $\Omega_{ij}$ tells us the probability of student $i$ picking student $j$ as best friends, $\Omega\Omega_{ij}$ informs the probability that student $i$ considers student $k$ as best friend who considers student $j$ as best friend. If transitivity is prevalent in classroom friendship, then we should observe that when $\Omega_{ij}$ is large then $\Omega\Omega_{ij}$ is also large. However, since we enforce the diagonal of $\Omega$ to be zero, we impose the same constraint on $\Omega\Omega$ and normalize each row to be a multinomial distribution. In short, the transitivity loss is computed as $T(X)=\lVert\Omega-(I-{\rm diag}(\Omega\Omega))^{-1}(\Omega\Omega-{\rm diag}(\Omega\Omega))\rVert_{2}^{2}$.
The loss function is a weighted sum of the three components: ${{\rm MSE}}+\kappa H(X)+\lambda T(X)={{\rm Bias}}^{2}+{{\rm Var}}+\kappa H(X)+\lambda T(X)$ as shown in Figure 1. So far, we have described the loss function for one classroom. The loss is summed over all classrooms.
Additional notes on loss function: When running the empirical analysis, we find it helpful to downweigh variance. The benefit of downweighing variance can be intuitively rationalized by considering network density. We defer the explanation to Appendix D.2 as it does not alter the main idea of why the loss function makes sense. The modified loss function can be written as ${{\rm Bias}}^{2}+\mu{{\rm Var}}+\kappa H(X)+\lambda T(X)$, where $(\mu,\kappa,\lambda)$ are hyperparameters. We can avoid cross-validation for tuning hyperparameters by examining whether $\Omega$ satisfies commonly known properties of social networks like gender homophily, presence of central nodes, and clustering. Such examination of $\Omega$’s interpretability can help pin down a range of reasonable values of $(\mu,\kappa,\lambda)$. We substantiate our claim in Appendix D.3.
4.2 Instrumental variable peer effect measurement
The existing peer effect literature largely adopts the following linear-in-means specification:
$\underbrace{y_{ics}}_{\text{Student i's 8th-grade cognitive test score}}=\beta\tilde{W}_{ics}+X_{ics}\gamma+\underbrace{\theta_{s}}_{\text{Fixed Effect (FE)}}+\underbrace{\mu_{cs}}_{\text{Random Effect (RE)}}+\epsilon_{ics}$ (2)
where $y_{ics}$ is cognitive ability test score of student $i$ from class $c$ school $s$, $\tilde{W}_{ics}=\frac{\sum_{j=1,j\neq i}^{N_{cs}}z_{jcs}}{N_{cs}-1}$ is student $i$’s middle school classmates’ average 6th-grade class quantile, $X_{ics}$ is a set of control variables for student $i$ such as the student’s 6th-grade class quantile, gender, race, and parent’s education level. $\theta_{s}$ and $\mu_{cs}$ are the school fixed effect and classroom random effect, respectively. We can rewrite linear-in-means model as the $y_{cs}=\beta W_{cs}z_{cs}+X_{cs}\gamma+\theta_{s}+\mu_{cs}+\epsilon_{ics}$ where
$W_{cs}=\begin{pmatrix}0&\frac{1}{N_{cs}-1}&\frac{1}{N_{cs}-1}&\dots&\dots&\frac{1}{N_{cs}-1}\\ \frac{1}{N_{cs}-1}&0&\frac{1}{N_{cs}-1}&\dots&\dots&\frac{1}{N_{cs}-1}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \frac{1}{N_{cs}-1}&\frac{1}{N_{cs}-1}&\frac{1}{N_{cs}-1}&\dots&\dots&0\end{pmatrix}.$
The linear-in-means model implicitly assumes that all students exert the same amount of peer influence on all their classmates as evidenced by the $W_{cs}$ matrix which has column sums equal to 1 for each column. This rules out the possibility of heterogeneous influence among all classmates and hence, no matter how we formulate classroom assignments, all students have the same total influence on all their classmates (sum up to 1) and there is no strategic scope to optimize classroom assignments. Eq(3) deviates from this setup.
Taking the $\Omega$ modeled by PeerNN as the true adjacency-probability matrix, we replace the linear-in-means specification with friendship-weighted specification, Eq(3), generating a more meaningful interpretation of $\beta$ for policymaker (See Appendix E for detail of linear-in-means model and its comparison to Eq(3)). However, we face an endogeneity issue with the $\Omega$ matrix which we illustrate with example C.4. Student $i$’s friendship selection, ${\Omega}_{ics}$, is correlated with some unobserved characteristics of student $i$ in $\epsilon_{ics}$.
$y_{ics}=\beta\tilde{\Omega}_{ics}+X_{ics}\gamma+\theta_{s}+\mu_{cs}+\epsilon_{ics}$ (3)
where $\tilde{\Omega}_{ics}={\sum_{j=1}^{N_{cs}}\Omega_{ijcs}z_{jcs}}$ is student $i$’s friends’ average 6th-grade ranking.
We resort to the classical instrumental variable (IV) approach to address the endogeneity. The instrument that we use is the average 6th-grade class quantile of middle school classmates, $\tilde{W}$ in Eq(2). A valid instrument needs to fulfill two properties: exclusion constraint and relevance constraint.
Exclusion constraint requires the instrument (average 6th-grade class rank for all classmates) to be uncorrelated with the unobserved confounder. This requirement is trivially satisfied since all classroom assignments are random conditional on school choice which is a covariate included in Eq(3). Hence, any confounder that is individual $i$-specific and predetermined before middle school should not correlate with the average characteristics of the classmates that are randomly assigned to him. Relevance constraint requires the instrument and the endogenous variable to be correlated. Both average 6th-grade class rank ($W_{cs}z_{cs}$) and friends’ weighted average 6th-grade class rank ($\Omega_{cs}z_{cs}$) are dependent on all classmates’ 6th-grade class rank ($z_{cs}$). Therefore, the relevance constraint is satisfied. We exemplify our explanations why both properties are satisfied with Example C.5.
We estimate Eq(3) using a two-stage method that resembles the classical two-stage least-square procedure:
  1. 1.
    Run OLS for reduced form equation (4) and get fitted values of $\tilde{\Omega}$.
    $\tilde{\Omega}_{ics}=\pi_{1}\underbrace{\frac{\sum_{j=1,j\neq i}^{N_{cs}}z_{jcs}}{N_{cs}-1}}_{\text{Classmates' average 6th-grade ranking}}+X_{ics}\delta+\phi_{s}+\eta_{ics}$ (4)
  2. 2.
    Substitute $\tilde{\Omega}$ in Eq(3) with the fitted value and run MLE assuming normally distributed $\mu$ and $\epsilon$.
We also implement other standard methods for IV and random effect (Balestra and Varadharajan-Krishnakumar, 1987; Baltagi, 1981; Han, 2016) to check the robustness of the regression results.
The two $\beta$s in Eq(2) and Eq(3) have different interpretations. For linear-in-means model, $\beta$ in Eq(2) is the marginal effect of middle school classmates’ average 6th-grade class quantile on the outcome; for friendship-weighted specification, $\beta$ in Eq(3) is the marginal effect of middle school friends’ weighted average 6th-grade class quantile on the outcome. [^1] Both marginal effect parameters (the two $\beta$s) reflect causality: the former comes from a randomized experiment whereas the latter uses IV to establish causality. The former has very little policy implication to classroom assignment since the linear-in-means model essentially restricts how much total peer effect each student can exert on other classmates (See detail in Appendix E). Our friendship-weighted specification relaxes this restriction: a popular student is allowed to exert more influence on all his classmates than another less popular student. This allows us to devise an average-maximizing class assignment if $\beta$ in Eq(3) is estimated to be statistically significant.
4.3 Genetic algorithm (GA) that optimizes class assignment
Since most of the schools have exactly two classrooms surveyed in CEPS, we assume the principal would like to optimize peer effect in these two classrooms, $c=\{1,2\}.$ [^2] Note that even though having more classrooms makes running GA more time-consuming, it makes the improvement of GA policy over random assignment more salient. This is because as we add in more classrooms, there are greater possibilities/configurations of classroom assignments for us to pull positively influencing student groups into one classroom and separate “disruptive peers” from those who are more likely to be influenced by them.
4.3.1 Genetic algorithm (GA)
In this section, we illustrate some specific features of the GA design that we use to optimize class assignment policy. We elaborate on the GA design that we use with Algorithm 1. Some of the referenced functions in Algorithm 1 are detailed in Appendix F. We summarize GA’s essence for those interested in the general concept of GA. Step 0: After a random initialization, GA evaluates the fitness of the initialized classroom assignment based on a fitness function (either Eq(5) or Eq(6)). Step 1: GA randomly swaps a pair of students to generate a new classroom assignment and evaluates its fitness. Step 2: GA repeats Step 1 with a new pair of students swapped each time. The number of repetitions is user-specified. Step 3: GA selects the best classroom assignment among those evaluated in Steps 0-2 and updates the optimal classroom assignment as that. Step 4: Repeat Steps 0-3 with the updated optimal classroom assignment as the initial assignment until some stopping criterion is satisfied. Note that this is a schematic summary of GA and many modifications are made when we apply GA. Readers should refer to Algorithm 1 for the exact implementation of our GA (and AFGA) design.
Algorithm 1 GA used to optimize class assignment policy
Input: $C$triangleright Set $C$ consists of all students the principal needs to divide into two classrooms
$N_{b}\leftarrow\sum_{s\in C}\mathbbm{1}_{\{s\text{ is boy}\}}$
$N_{g}\leftarrow\sum_{s\in C}\mathbbm{1}_{\{s\text{ is girl}\}}$
$N\leftarrow N_{b}+N_{g}$triangleright Alternatively, $N=\lvert C\rvert$
$j=\operatorname*{\arg\!\min}_{j}\{N_{j}\}_{j\in\{b,g\}}$
do
    $v\leftarrow$ Randomly draw $\lfloor{\frac{\lvert C\rvert}{2}}\rfloor$ students from $C$ without replacement and uniform probability
    $C_{1}\leftarrow C[v]$
    $C_{1}\leftarrow C\backslash C_{1}$triangleright This initialization ensures that $-1\leq\lvert C_{1}\rvert-\lvert C_{2}\rvert\leq 1$
while check$(C_{1},C_{2},N_{b},N_{g},j)=0$triangleright See Algorithm 2 in Appendix F
L = 150
M = 100
$PH={\rm matrix}(0,L,N)$triangleright Record all the chosen class assignment policies
$FSH={\rm rep}(0,L)$triangleright Record all fitness scores of the chosen policies
for $l$ in $1:L$ do
    $b\leftarrow B\sim{\rm Bernoulli}(0.05)$
    if b = 1 then $\triangleright$ Mutation starts
         do
             $t_{1}\leftarrow T_{1}\sim{\rm Multinomial}(n=1,{\rm prob}={\rm rep}(\frac{1}{\lvert C_{1}\rvert},C_{1}))$
             $t_{B}\leftarrow T_{B}\sim{\rm Multinomial}(n=1,{\rm prob}={\rm rep}(\frac{1}{\lvert C_{2}\rvert},C_{2}))$
             $CP\leftarrow(C_{1}[t_{1}],C_{2}[t_{2}])$
             $C_{1}^{\prime},C_{2}^{\prime}\leftarrow{\rm swap}(C_{1},C_{2},CP)$triangleright See Algorithm 3 in Appendix F
         while check$(C_{1}^{\prime},C_{2}^{\prime},N_{b},N_{g},j)=0$triangleright Mutation ends
    $\triangleright$ If not mutation, then crossover
         $FS={\rm rep}(0,M)$
         $v_{1}\leftarrow V_{1}\sim{\rm Multinomial}(n=M,{\rm prob}={\rm rep}(\frac{1}{\lvert C_{1}\rvert},C_{1}))$
         $v_{2}\leftarrow V_{2}\sim{\rm Multinomial}(n=M,{\rm prob}={\rm rep}(\frac{1}{\lvert C_{2}\rvert},C_{2}))$
         $CS\leftarrow(C_{1}[v_{1}],C_{2}[v_{2}])$triangleright $CS$ denotes M candidate swaps
         for $m\in 1:M$ do $\triangleright$ Crossover loop
             $(C_{1}^{\prime},C_{2}^{\prime})={\rm swap}(C_{1},C_{2},CS_{m})$triangleright $CS_{m}$ is the $m$th candidate swap
             if check$(C_{1}^{\prime},C_{2}^{\prime},N_{b},N_{g},j)=1$ then
                 $FS[m]\leftarrow{\rm fit}(C_{1}^{\prime},C_{2}^{\prime})$triangleright ‘fit’ is either Eq(5) or Eq(6)
             else
                 $FS[m]\leftarrow-\infty$
             end if
         end for
         $bp\leftarrow\operatorname*{\arg\!\max}_{m}FS[m]$
         if $FS[bp]>FSH[l-1]$ then $\triangleright$ Better than last round chosen policy
             $CP\leftarrow CS_{bp}$
         else
             $CP\leftarrow({\rm NULL},{\rm NULL})$
         end $\triangleright$ Crossover ends
    end if
    $(C_{1},C_{2})\leftarrow{\rm swap}(C_{1},C_{2},CP)$
    $PH[l,]=(C_{1},C_{2})$
    $FSH[l]={\rm fit}(C_{1},C_{2})$
end for
$bp\leftarrow\operatorname*{\arg\!\max}_{l}FSH[l]$
$C_{1},C_{2}\leftarrow PH[bp,]$
We use a selection-mutation-crossover-termination four-step design. One special feature of our GA design is the incorporation of gender-ratio constraint: a swap that violates this gender ratio constraint will not be permitted. Given that all schools in our sample have close to a 50% gender ratio, multitudes of candidate policies that satisfy the gender ratio requirement are guaranteed to exist. Additionally, we constrain that the classrooms’ sizes differ by at most 1, ensuring the classrooms are of comparable sizes. See Appendix F for a detailed description of the institutional constraint and the parameter values that we set for GA.
4.3.2 Fitness function
The most natural fitness function for a principal is the mean predicted peer effect for all students, $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}$, where $\tilde{\Omega}_{c}=\Omega_{c}z_{c}$ is a vector of each student’s friends’ 6th-grade average class quantile. This fitness function specification potentially causes algorithmic fairness problems. GA may maximize the average at the expense of certain small groups of students’ educational outcomes. We actively combat this problem by penalizing both the standard errors of peer effect within classrooms ${\rm SE}_{\{\beta\tilde{\Omega}_{1}\}}$ and ${\rm SE}_{\{\beta\tilde{\Omega}_{2}\}}$ as well as standard error of peer effect across classrooms ${\rm SE}_{\{\beta\tilde{\Omega}\}}$, where $\tilde{\Omega}=(\tilde{\Omega}_{1},\tilde{\Omega}_{2})$. The two fitness functions are given as follows:
${\rm fit}_{{\rm GA}}(C_{1},C_{2})=$ $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}$ (5)
${\rm fit}_{{\rm AFGA}}(C_{1},C_{2})=$ $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}-\sum_{c=\{1,2\}}\phi{\rm SE}_{\{\beta\tilde{\Omega}_{c}\}}-\rho{\rm SE}_{\{\beta\tilde{\Omega}\}}.$ (6)
where $C_{1}$ and $C_{2}$ denote the set of students assigned to the two classrooms, respectively. Note that $(C_{1},C_{2})$ fully pins down the class assignment policy, $\pi$. The penalties are weighted by $\phi$ and $\rho$.
5 Results
5.1 Friendship formation pattern
Figure 2: Heat map for classroom 1’s adjacency-probability matrix $\Omega$. The $(i,j)$-th entry, $\Omega_{ij}$, denotes the probability of student $i$ considering student $j$ as best friend. Students are sorted by gender. In Appendix G, we show that the friendship formation patterns reflected in this figure is also true in other classrooms.
PeerNN can capture the most salient features in the classroom network. First, it predicts extremely strong gender homophily. We reorder all students by gender in the first classroom in our dataset and plot a heat map of $\Omega$, the adjacency-probability matrix (See Figure 2). Two dark diagonal blocks indicate that students are more likely to make friends of the same gender as themselves. Second, PeerNN finds some popular individual students, a common observation in social networks. Many social networks exhibit centrality property where certain nodes are particularly well-connected with other nodes. Third, PeerNN is, to some extent, able to capture clustering within gender groups. We observe from Figure 2 that each student has heterogeneous popularity across classmates of the same gender group, a necessary condition implied by clustering.
We compare PeerNN with the conventional linear-in-means model. CEPS asks students to provide ARD responses about their friends in terms of ten distinct traits. We use $\Omega$ to predict friendship formation for all students and compare our prediction with the ARD responses. How we compute prediction error is detailed in Algorithm 4 in Appendix G. In Figure 3, we present violin plots of prediction error. For half of the ten traits including gender, local/nonlocal, smoke or drink, our model predicts better for all 1000 rounds. PeerNN also predicts better than the linear-in-means model for the rest of the traits except for ‘hardworking’. It could be due to the fact students’ own evaluation of whether they are hardworking systematically differs from their friends’ evaluation of them. Nevertheless, Overall, our model beats the linear-in-means model in terms of out-of-sample prediction by a huge margin.
Figure 3: PeerNN prediction performance full version. Blue and yellow color distributions represent the empirical densities of PeerNN and linear-in-means models’ prediction errors, respectively. PeerNN outperforms the linear-in-means model in nine out of ten traits. The only trait that PeerNN performs worse than the linear-in-means model. It is likely due to how students evaluate whether they are hardworking or not may systematically differ from how they evaluate whether their friends are hardworking or not. Also, note that the scale for the ‘hardworking’ trait is much smaller than some of the other traits such as gender and local.
We present all ten violin plots in Figure 3. CEPS asks students to provide information about their friends in terms of ten distinct traits. We randomly draw without replacement $B_{i}$ friends for a student based on the multinomial distribution $\Omega_{i}$, the $ith$ row of $\Omega$, predict each of the traits, and record the prediction error for each trait. We repeat this step 1000 times for each student. For half of the ten traits including gender, local/nonlocal, smoke or drink, our model predicts better for all 1000 rounds. For the rest of the traits, our model prediction is better than the linear-in-means model most of the time except for the ‘hardworking’ trait. Overall, our model beats the linear-in-means model in terms of out-of-sample prediction by a huge margin.
5.2 Positive peer effect from ‘good students’
We estimate Eq(3) and confirm the presence of positive peer effect from ‘good students’ (students who did well in elementary school). In particular, if student $i$’s friends’ weighted average 6th-grade class quantile improves by 10%, his 8th-grade cognitive ability test score improves by 0.1082 on a scale of 5, or equivalently, 13% of standard deviation. To put it more concretely, the average cognitive test score is 0.3608 which lands the student at the 43rd quantile of the score distribution (the score distribution is left-skewed). An 0.1082 increase in cognitive test score improves the student’s ranking from 43rd quantile to 51st quantile. In Table 2, we show four specifications’ results.
  • Linear-in-means model without classroom random effect
  • Linear-in-means model with classroom random effect
  • Instrumental variable without classroom random effect
  • Instrumental variable with classroom random effect
Table 2: The Peer Effects of Sixth Grade Rank on Cognitive Test Score
Linear-in-means Linear-in-means IV IV
Peer’s Rank 0.973∗∗∗ 0.980∗∗ 1.075∗∗∗ 1.082∗∗
(0.244) (0.462) (0.269) (0.510)
Own Rank 1.019∗∗∗ 1.022∗∗∗ 0.995∗∗∗ 0.997∗∗∗
(0.038) (0.039) (0.038) (0.038)
Age -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗
(0.001) (0.001) (0.001) (0.001)
Sex -0.010 -0.010 -0.011 -0.010
(0.017) (0.017) (0.017) (0.017)
Father’s education 0.020∗∗∗ 0.017∗∗∗ 0.020∗∗∗ 0.018∗∗∗
(0.006) (0.006) (0.006) (0.006)
Mother’s education 0.006 0.006 0.005 0.005
(0.006) (0.006) (0.006) (0.006)
Ethnic nationality 0.013 0.022 0.012 0.021
(0.040) (0.040) (0.040) (0.040)
Number of observations 5,860 5,860 5,860 5,860
Adjusted $R^{2}$ 0.382 0.382
- Log Likelihood 5,845.390 5,845.291
Results from first stage
Peer’s Rank 0.905∗∗∗ 0.905∗∗∗
(0.015) (0.015)
Number of observations 5,860 5,860
Adjusted $R^{2}$ 0.755 0.755
School FE YES YES YES YES
Class RE NO YES NO YES
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
We also implement more standard methods (Baltagi, 1981; Balestra and Varadharajan-Krishnakumar, 1987) for IV with random effect specification to estimate Eq(3). The estimated coefficients for peer effect are 0.975 (0.268) and 1.075 (0.270), respectively. The numbers in the parenthesis are standard errors. The regression results for all four specifications and all estimation methods are very close to each other, demonstrating robustness. See source code in supplemental material for details.
The regression results opens up the strategic scope of classroom assignments. Making good students the central nodes of the social network increases the entire classroom’s peer effect. On the other hand, if the estimated $\hat{\beta}$ were significantly negative, we would aim to make good students peripheral nodes. The sign of the estimated peer effect determines two opposing sets of policies.
5.3 Optimal classroom assignment policy design
Figure 4: We plot $Q$ matrices under raw class assignment, GA assignment, and AFGA assignment. The indices of these two classrooms are 5 and 6. Two columns and rows of dark grids in Raw Q2 indicate that in classroom 6, two disruptive peers negatively influence the entire classroom to a large extent. The extremely dark grids in GA Q2 indicate that GA gives up on a few students and lets them form a clique that is almost exclusively- and negatively-influencing those in the clique. AFGA has a higher average peer effect than GA as shown by the lighter color (on average) of AF-GA plots than Raw plots. AFGA also does not suffer from the extreme inequity problem as in GA Q2. Appendix H shows that the same patterns persist throughout various classrooms/schools.
To assist visualizing the class assignment results for a single classroom, we define the demeaned peer effect $q_{ij}$ for each pair of student $\{i,j\}$, $q_{ij}=\Omega_{ij}\tilde{z}_{j}+\Omega_{ji}\tilde{z}_{i}$ where $\tilde{z}_{i}=z_{i}-\frac{1}{N}\sum_{j=1}^{N}z_{j}$. We collect all paired peer effect, $q_{ij}$, into a matrix $Q$, and since $q_{ij}=q_{ji}$ by definition, $Q$ is an $(N\times N)$ symmetric matrix. Technically, $\beta Q$ is each pair’s predicted demeaned peer effect, but since $\beta>0$ is a constant, we omit it. In Figure 4, we illustrate $Q$ as a heat map, with the term ‘raw’ assignment referring to the initial classroom assignment policy randomly generated by the school principal.
Under the raw assignment, many students are negatively influenced by two disruptive peers as indicated by the two darker rows and columns of the Raw Q2 plot in Figure 4. GA separates many pairs of students who suffer from such large negative peer effects by ensuring that these two disruptive peers are popular among fewer students as evidenced by the reduced number of dark grids in GA Q1 and GA Q2 relative to Raw Q1 and Raw Q2. Moreover, GA Q2 shows that the entire classroom mostly consists of positively influencing friend pairs. Consequently, students’ average predicted peer effect, Eq(5), increases by 1.9% under GA assignment policy. Unfortunately, maximizing the average peer effect inadvertently leads to severe negative peer effects for some students. In GA Q2, many students face extremely negative peer effects.
[…]
We deploy AFGA to address this inequity issue. AF-GA Q1 and AF-GA Q2 in Figure 4 show a much more equitable educational outcome. None of the students suffer from extremely negative peer effects as in the GA Q2. However, though it is not clear in Figure 4, the average peer effect for AFGA (compared to raw assignment, the improvement is 1.2%) is lower than GA (1.9%), which is expected since the objective function of AFGA Eq(6) does take into account equity. If the school principal believes this particular policy designed by AFGA overvalues equity and sacrifices efficiency too much, he can decrease $\phi$ and $\rho$ for a more efficient outcome, however, the outcome will be less equitable. Note that the school principal can predict peer effect, $\tilde{\Omega}$, for all students before implementing the policy, hence he can always choose $\phi$ and $\rho$ based on his preference over efficiency versus equity.
In Figure 5, we compare the equity of all three policies more directly. Raw class assignment policy generates a trimodal distribution of realized peer effect. The leftmost mode is due to the strong influence of disruptive peers. GA essentially breaks the friend circle that generates the leftmost mode. However, in the process of pursuing the highest average, it ignores the peer effect for a small number of students, generating extremely inequitable educational outcomes as evidenced by the few left-tail outliers. AFGA takes into account the standard deviation of the entire distribution, outputting a much fairer policy rule.
6 Conclusion
This study constructs a micro-founded model for predicting friendship formation and employs a novel and interpretable neural network architecture called PeerNN to estimate this model. Leveraging the predictions generated by PeerNN, we consistently estimate peer effects using a linear-in-means instrument. By combining the results of friendship formation prediction and peer effect parameter estimation, we simulate counterfactual peer effects for all students. Our work then designs an algorithmic fair genetic algorithm, outputting a class assignment policy that enhances average peer effects while maintaining fairness.
Broader impact: This three-step empirical framework is extendable to other group assignment challenges where peer effects play a pivotal role. Some of the applicable contexts, in which peer effects have been empirically proven crucial and group assignments are common practices, are workplace teams (Anderson et al., 2008; Bandiera et al., 2010; Mas and Moretti, 2009), healthcare interventions (Pianese and Belfiore, 2021; Pagkas-Bather et al., 2020) and community programs (Valente et al., 2003; Yu et al., 2024; Cero et al., 2024). Similar to classroom assignments, navigating the trade-off between efficiency and equity in these situations can present moral dilemmas.
Limitations and future work: The scope of this paper does not include the theoretical properties of the three-step empirical framework. Some of the future research directions are (1) characterizing the conditions under which the friendship formation model is identifiable, (2) proving consistency of the IV approach when $\Omega$ is not taken as the ground truth, and (3) showing the error bound of GA as the number of search goes to infinity.
Appendix A Further literature review
A.1 Comparison between our model and other causal inference methods under networked interference
Ma and Tresp (2021); Jiang and Sun (2022); Huang et al. (2023) all adapt GNN for causal inference. The core problem in treatment effect estimation is that the treated group and control group have imbalanced features, both their individual feature as well as their neighborhood information.
For example, students who enroll themselves in a tutoring program are inherently different (possibly more self-motivating) from those who don’t. At the same time, these tutoring program participants may have friends who tend to do well academically. Hence, if we observe tutoring program participants on average do better than other students, we cannot make the casual statement that attending the tutoring program improves students’ grades. Their better performance could be due to their self-motivation or more helpful friends.
To resolve these two imbalanced data problems, all three papers encode their raw data into latent feature space. They formalize mathematically the concept of balancing the distributions of latent features for treated and control groups as part of the loss function. Ma and Tresp (2021); Huang et al. (2023) both directly penalize some distance measures of the two distributions (latent feature distribution of the treated group and that of the control group). Jiang and Sun (2022), in contrast, trains a discriminator that learns about the probability of treatment based on those latent features and eventually outputs a set of latent features such that the discriminator can no longer distinguish the treatment probability between the treated group and control group. For an imbalanced network problem, all three papers also collapse all neighbors’ treatment information into a scalar, percentage of neighbors being treated. After the transformation of neighbor information, they address the network imbalance problems just like how they resolve the treatment imbalance.
Additionally, Huang et al. (2023) controls for unobserved confounding by adding the treatment prediction loss to its loss function. The weights for balancing the distributions and treatment prediction are tuning parameters in their case.
The challenge faced by those three papers is rooted in nonrandom treatment. In our case, the class assignment itself is random for 5860 students and 139 classes. Nevertheless, causal inference is not a straightforward linear regression because friendship formation is not random. We exploit the random assignment to generate instrumental variable for friends’ weighted average 6th-grade class rank. By standard two-stage least-square estimation, we can consistently estimate the parameter of interest $\beta$ in Equation (3). The instrumental variable identification strategy solves both the ‘treatment’ imbalance (we can perceive friends’ weighted average 6th-grade class quantile as continuous treatment) and unobserved confounding problems.
Next, we compare the policy optimization method of our paper and Ma and Tresp (2021). Ma and Tresp (2021) build a neural network to solve for optimal policy. This is doable in their case as the network is predetermined and treatment assignment does not change the network. In contrast, our class assignment changes the network structure. For example, if students $i$ and $j$ are not in the same class, we assume that there is zero probability of being friends, on the other hand, when they are in the same classroom, there is always a nonzero probability of them being friends due to the use of softmax function. Network structure is assignment-dependent in our case. This means that if we were to build a neural network to solve for optimal policy, we would need to update the network probability matrix for each backpropagation. It is computationally infeasible. Therefore, we resort to a popular heuristic combinatorial optimization method: genetic algorithm.
A.2 Evolutionary algorithm
We highlight the similarities, differences, and relevant applications of Yu et al. (2023) to our AFGA design, a fairly new paper in the evolutionary algorithm literature. While doing so, we include other relevant papers as well.
  • While evolutionary learning is popular in feature selection, task scheduling, and game theory (Yu et al., 2023; Yang et al., 2022; Li and Li, 2020), we bring a subset of evolutionary learning (GA) to a new domain: network-dependent group assignment.
  • Just like Yu et al. (2023), we treat the optimization problem as a combinatorial one in contrast to the popular DARTS method (Liu et al., 2018) which reformulates the optimization problem as a continuous one.
  • Conditionally on a random draw, the mutation stage of AFGA updates the class assignment policy “deterministically”, same as (1,1)-CELS instantiation in Yu et al. (2023).
  • Yu et al. (2023)’s (1+1)-CELS mutates parents that are hard to beat with higher probability. This concept can be adapted by our method by replacing ${\rm Multinomial}(n=1,{\rm prob}={\rm rep}(\frac{1}{\lvert C_{1}\rvert},C_{1}))$ with ${\rm Multinomial}(n=1,{\rm prob}=1/((\sum 1/\Omega z)\Omega z))$ in Algorithm 3. The modification is similar to (1+1)-CELS as we are also encouraging mutations for undesirable policy configuration (local optima in Yu et al. (2023), negatively-influencing pairs in our paper).
  • Similar to the last point, we can adapt population-based search (He and Yao, 2002; Mallipeddi and Suganthan, 2008; Tang et al., 2016; Yu et al., 2023) to our setup. At the crossover stage, we randomly select pairs of students to swap. We can assign a higher probability of being swapped to students from classrooms with low average peer effects or large variance of peer effects.
Appendix B Friendship questionnaire
The National Survey Research Center at Renmin University of China (NSRC) provides CEPS and all information related to the dataset. The questionnaire can be found at http://ceps.ruc.edu.cn/Enlish/dfiles/11184/14391213173188.pdf (English version) and http://ceps.ruc.edu.cn/__local/C/67/E7/F7BA6FBDFBC3C8B808AA0F6E0FA_905BECDA_7A84E.pdf?e=.pdf (Chinese version). In Figure 6, we show a screenshot of the ARD questions from the 7th-grade survey. More ARD questions are not included in the screenshot, interested readers can use the aforementioned URLs to look at the questions. The data that we use is from the 8th grade survey, however, the ARD questions from the 8th grade survey are the same as the ARD questions from the 7th grade. We provide the screen of the 7th-grade survey as the English translation version is only available for the 7th-grade survey, not the 8th-grade survey.
Figure 6: The friendship data is aggregated relational. The identification of network formation parameters with such data is studied by Breza et al. (2020). We focus on friendship formation prediction instead of inference on the preference parameters. For section C21, there are nine questions in total. We do not include all the questions in section C21 in the figure. We will use students’ responses to these questions ($A_{f}$) as the response variable for PeerNN.
Appendix C Examples
Example C.1.
A typical mainstream problem that fits current literature and our research context would be: if a new student joins a classroom, given the current network structure in the classroom, all classmates’ characteristics, and the new student’s characteristics, who is most likely to befriend the new student (i.e. predicting future link formation). This is a factual prediction problem as we can observe the friendship decisions made by the new student later on. Our work solves a different problem. In our work, we are interested in the following problem: given the current network structures of some 8th-grade classrooms and all students’ 6th-grade characteristics, how will the network structures change if we assign students differently two years ago at the start of these students’ 7th-grade. This is a counterfactual prediction problem. We will never observe the counterfactual scenario that we are trying to predict.
In the former question, one may rely more on graph-based measures such as node degree to make a prediction; in the latter, we use content-based measures exclusively to predict friendship. See Daud et al. (2020) for details on graph-based and content-based measures. Since the latter type of questions including our research question can only use content-based measures, there is arguably a greater need for learning about the underlying friendship preference function than the former type of questions.
More importantly, the former question is a factual prediction problem, whereas the latter is a counterfactual prediction problem. While researchers can observe how a new student forms friendships in a classroom, they never observe the same group of students being assigned to different classrooms twice and observe two sets of networks independently. Therefore, many of the techniques that we use for prediction (e.g. any method that uses existing network structure such as using GNN to aggregate neighbor information) cannot be applied or have to be applied differently to the latter question.
With example C.1, we argue that a micro-founded approach supported by economic utility theory is particularly suitable for our research question.
Example C.2.
Instead of using latent representations $\sigma(X)$, we are going to use raw data features, $X$, to illustrate the friendship demand concept. We find this simplified illustration (using $X$ instead of $\sigma(X)$) more intuitive, but it still captures the essence of the microfoundation of the model. Stage 0 (latent feature encoder) is not of the essence of the four stages in the micro-founded friendship prediction model. One can drop Stage 0 completely, it does not affect the training result much. We add Stage 0 to improve the flexibility of the neural network.
Use two predictors: gender and whether a student did well in 6th grade to predict friendship formation. We have the data matrix (Table 4) for a classroom consisting of five students.
gender did well in 6th-grade
Adam 1 1
Ben 1 1
Cam 1 0
Debbie 0 1
Emily 0 0
Table 3: Without encoder, $\sigma(X)=X$ contains students’ predetermined characteristics: gender and whether the student did well in 6th-grade
gender did well in 6th-grade
Adam 1 0.5
Ben 1 0.5
Cam 0.5 -0.5
Debbie -1 0.5
Emily -0.5 -0.5
Table 4: Preference parameter $\delta$ indicates students’ preference for making friends with classmates with certain characteristics
Since both input variables are binary, there are four types of students. (1) boy who did well, (2) boy who did not do well, (3) girl who did well, and (4) girl who did not do well. Assume that the four types of students have the following preferences:
  • Boys who did well in elementary school prefer to make friends with boys over girls and they prefer to make friends with those who did well in elementary school.
  • Boys who did not do well in elementary school prefer to make friends with boys over girls and they prefer to make friends with those who did not do well in elementary school.
  • Girls who did well in elementary school prefer to make friends with girls over boys and they prefer to make friends with those who did well in elementary school.
  • Girls who did not do well in elementary school prefer to make friends with girls over boys and they prefer to make friends with those who did not do well in elementary school.
A possible $\delta$ matrix to describe such a preference pattern is Table 4. Note that $\delta_{i}$ is a function of $X_{i}$ and since Adam and Ben share the same $X$ values, their preference parameters are identical.
Example C.3.
We compute the outer product of $\delta$ and $\sigma$ from example C.2 to compute the linearized propensity score $\Upsilon_{ij}$.
$\begin{pmatrix}1&0.5\\ 1&0.5\\ 0.5&-0.5\\ -1&0.5\\ -0.5&0.5\end{pmatrix}\begin{pmatrix}1&1&1&0&0\\ 1&1&0&1&0\end{pmatrix}=\begin{pmatrix}1.5&1.5&1&0.5&0\\ 1.5&1.5&1&0.5&0\\ 0&0&0.5&-0.5&0\\ -0.5&-0.5&-1&0.5&0\\ 0&0&-0.5&0.5&0\\ \end{pmatrix}=\Upsilon$
We then set the diagonal entries of $\Upsilon$ to be negative infinity such that the probability of a student making friends with himself is 0. After that, we apply a softmax function to each row of $\Upsilon$.
$\begin{pmatrix}1.5&1.5&1&0.5&0\\ 1.5&1.5&1&0.5&0\\ 0&0&0.5&-0.5&0\\ -0.5&-0.5&-1&0.5&0\\ 0&0&-0.5&0.5&0\\ \end{pmatrix}\overset{\Upsilon_{ii}=-\infty}{\implies}$ $\begin{pmatrix}-\infty&1.5&1&0.5&0\\ 1.5&-\infty&1&0.5&0\\ 0&0&-\infty&-0.5&0\\ -0.5&-0.5&-1&-\infty&0\\ 0&0&-0.5&0.5&-\infty\\ \end{pmatrix}$
[…]
Example C.4.
Students may choose to make friends with those of similar IQ as themselves. Higher IQ also has a positive impact on students’ 8th-grade cognitive ability test scores. We do not observe IQ and hence it is part of the error term. Then, if we run regression based on Eq(3), we may overestimate peer effect $\beta$ since we inevitably attribute some of the impact of IQ to peer effect.
Example C.5.
We continue with example C.4 where IQ is an endogenous variable.
Exclusion constraint requires the instrument (average 6th-grade class rank for all classmates) to be uncorrelated with the unobserved confounder IQ. This requirement is trivially satisfied since all classroom assignments are random conditional on school choice which is a covariate included in Eq(3).
Relevance constraint requires the instrument and the endogenous variable to be correlated. Both average 6th-grade class rank ($W_{cs}z_{cs}$) and friends’ weighted average 6th-grade class rank ($\Omega_{cs}z_{cs}$) are dependent on all classmates’ 6th-grade class rank ($z_{cs}$). Therefore, the relevance constraint is satisfied.
Example C.6.
A principal needs to divide 100 students into two classrooms, each consisting of 50 students. There are $\frac{100!}{50!}$ ways to assign students to classrooms.
Appendix D Explanation for the loss function
D.1 Fitted value’s MSE
CEPS provides aggregated relational data. We design Stage 4 of PeerNN in accordance with this data feature. If researchers have network linkage data, this stage can be simplified.
Given a $(N\times N)$ matrix $\Omega$ whose rows are multinomial distributions, $\Omega_{ij}$ is the probability of students $i$ considering student $j$ as a friend. The diagonal of $\Omega$ is 0. Each row of $\Omega$ sums to 1. If we knew the linkage data, we could do a classical maximum likelihood estimation. In the case of aggregate relational data, we can randomly draw friends for all students based on $\Omega$ without replacement. However, this gradient-estimation-based method can be computationally intensive as $\Omega$ is a large probability matrix. We make two compromises to reduce the computation complexity. Nevertheless, the out-of-sample performance of PeerNN beats the linear-in-means model by a long shot, hence, we feel comfortable with making those compromises. Future research may build more reasonable micro-founded models to predict friendship formation.
D.1.1 Compromise 1
We let students draw best friends with replacement. In reality, when students answer questions in Figure 6, they are unlikely to put down the same friend twice. However, a random draw with replacement does not rule out such a possibility. Nevertheless, this compromise exponentially decreases the computational burden. We show that with the following scenario.
Assume there is a class consisting of $N$ students, and all students report five best friends. Let $V_{i}$ be the randomly drawn best friends vector for student $i$ without replacement. $V_{i}$ is of length $N$, the ith entry of $V_{i}$, $V_{ii}=0$ because $\Omega_{ii}=0$, and five of the entries of $V_{i}$ is 1 while the rest of the entries are 0. We break MSE into bias square and variance. In order to compute the bias square, we need to compute the expectation of $V_{i}$. For the first friend, computing expectation is straightforward $E[V_{i(1)}]=\Omega_{i}$. Computing for the expectation of the first two friends is exponentially more computationally burdensome.
$E[V_{i(1)}+V_{i(2)}]$ $=E[V_{i(1)}+E[V_{i(2)}|V_{i(1)}]]$
$=\Omega_{i}+E[\frac{\Omega_{i}-V_{i(1)}\circ\Omega_{i}}{1-V_{i(1)}\cdot\Omega_{i}}]$
$=\Omega_{i}+\sum_{j\neq i}^{N}\Omega_{ij}\frac{\Omega_{i}-e_{j}\circ\Omega_{i}}{1-e_{j}\cdot\Omega_{i}}$
$e_{j}$ is a vector of zeros except for the jth entry being 1. $\circ$ and $\cdot$ denote elementwise product and dot product, respectively. Using the same method as in the chain of equality (law of iterated expectation), we can extend the result to the first three draws of best friends without replacement.
$E[V_{i(1)}+V_{i(2)}+V_{i(3)}]=\Omega_{i}+\sum_{j\neq i}^{N}\Omega_{ij}\frac{\Omega_{i}-e_{j}\circ\Omega_{i}}{1-e_{j}\cdot\Omega_{i}}+\sum_{k\notin{i,j}}^{N}\sum_{j\neq i}^{N}\Omega_{ik}\Omega_{ij}\frac{\Omega_{i}-e_{j}\circ\Omega_{i}-e_{k}\circ\Omega_{i}}{1-e_{j}\cdot\Omega_{i}-e_{k}\cdot\Omega_{i}}$
It is clear that as the number of drawn best friends that we consider increases, the computational complexity increases exponentially. For the case that we consider where all students report 5 friends, the computational complexity for computing the expectation of $V_{i}$ only is $O(N^{5})$. For each student, the computational complexity is $O(N^{4})$ and there are $N$ students that we need to compute bias for. Moreover, since CEPS provides aggregate relational data only, our loss function will be even more complex than computing the expectation of $V_{i}$ only. As such, we make the ‘with replacement’ compromise. We will show that with compromises 1 and 2, the complexity of the MSE component of the loss function is reduced to $O(N^{3})$.
D.1.2 Compromise 2
We use five linear functions to approximate the correspondence between drawn friends’ characteristics and students’ evaluation of their friends. The survey design of CEPS forces surveyed students to report their responses in $\{1,2,3\}$ as depicted in Figure 6. The prediction that we make should match students’ responses about their best friends.
Given $\Omega$ and students’ self-evaluation $A_{s}$, we can randomly draw friends for students and predict, for example, how many of their friends study hard based on randomly drawn friendship vector $V_{i}$ and the column in $A_{s}$ which corresponds to an indicator of whether a student studies hard or not. The product of $V_{i}A_{s}$ informs us how many of the best friends whom student $i$ reports exhibit a trait such as studying hard. Note that $V_{i}A_{sq}\in\{0,1\dots,B_{i}\}$ where $B_{i}$ is the number of reported best friends up to 5. Hence, there needs to be a mapping from $(V_{i}A_{s},B_{i})$ to students’ evaluation of their best friends, $A_{fi}$, we dropped the subscript for question $q$ since the mapping is the same for all questions. $A_{s}$ and $A_{fi}$ are both $(N\times 10)$ matrices. Table 5 documents the correspondence or multi-valued function, $\mathcal{G}:\mathcal{B}\times\mathcal{V}\times\mathcal{A}\to\{1,2,3\}$.
$V_{i}\cdot A_{sq}=0$ $V_{i}\cdot A_{sq}=1$ $V_{i}\cdot A_{sq}=2$ $V_{i}\cdot A_{sq}=3$ $V_{i}\cdot A_{sq}=4$ $V_{i}\cdot A_{sq}=5$
$B_{i}=1$$\{1\}$$\{2,3\}$$\times$ $\times$ $\times$ $\times$
$B_{i}=2$$\{1\}$$\{2\}$$\{2,3\}$$\times$ $\times$ $\times$
$B_{i}=3$$\{1\}$$\{2\}$$\{2,3\}$$\{3\}$$\times$ $\times$
$B_{i}=4$$\{1\}$$\{2\}$$\{2\}$$\{3\}$$\{3\}$$\times$
$B_{i}=5$$\{1\}$$\{2\}$$\{2\}$$\{3\}$$\{3\}$$\{3\}$
Table 5: Students’ responses to friendship questionnaire may cause ambiguity. Given $B_{i}$ and $V_{i}\cdot A_{sq}$, we cannot uniquely pin down students’ responses to the friendship questionnaire.
There are two problems with $\mathcal{G}$. One, its output can be multi-valued and we cannot uniquely pin down students’ evaluation of their friends even if we know who their friends are. Two, it is nonlinear and renders ${\rm Bias}^{2}$ and ${\rm Var}$ to have no closed forms. We approximate $\mathcal{G}$ with five linear functions. The five linear functions, named $g$, are depicted in Figure 7. $g_{B_{i}}(V_{i},A_{s})=a_{B_{i}}V_{i}\cdot A_{s}+b_{B_{i}}$ is linear in $V_{i}$, the only random variable in function $g$. We collect the intercept and slope of these five linear functions in Table 6.
Figure 7: Five linear functions $g$ approximate the correspondence $\mathcal{G}$. The dots are the output values of $\mathcal{G}$ (i.e. values in Table 6). The lines are the least square linear best fit of those dots. We need $g$ to be linear in $V_{i}$ and hence, allow the intercept of slope to vary with $B_{i}$.
$B_{i}$ 1 2 3 4 5
$b_{B_{i}}$ 1. 1.090 1.154 1.2 1.333
$a_{B_{i}}$ 1.5 0.727 0.654 0.5 0.4
Table 6: Intercept and slope coefficients $g$.
D.1.3 Decomposition of MSE
We can work out the decomposition as a bias-variance formula.
$E[\overbrace{(g(V_{i}A_{s},B_{i})}^{Q\times 1\text{vector}}-A_{fi})^{\top}(g(V_{i}A_{s},B_{i})-\overbrace{A_{fi}}^{Q\times 1\text{vector}})]$
$=$ $E[(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})]+E[g(V_{i}A_{s},B_{i})]-A_{fi})^{\top}$
$(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})]+E[g(V_{i}A_{s},B_{i})]-A_{fi})]$
$=$ $\underbrace{(E[g(V_{i}A_{s},B_{i})]-A_{fi})^{\top}(E[g(V_{i}A_{s},B_{i})]-A_{fi})}_{\text{bias square}}+$
$\underbrace{E[(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])^{\top}(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])]}_{\text{`variance'}}+$
$E[2\underbrace{(E[g(V_{i}A_{s},B_{i})]-A_{fi})^{\top}}_{\text{constant}}\underbrace{(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])}_{\text{zero expectation}}]$
Next, we show how to relate the ‘variance’ term to the variance of $g(V_{i}A_{s},B_{i})$.
$E[(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])^{\top}(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])]$
$=$ $E[tr((g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])^{\top}(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})]))]$
$=$ $E[tr((g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])^{\top})]$
$=$ $tr(E[(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])(g(V_{i}A_{s},B_{i})-E[g(V_{i}A_{s},B_{i})])^{\top}])$
$=$ $tr({\rm Var}(g(V_{i}A_{s},B_{i})))$
D.1.4 Closed-form expressions for ${\rm Bias}^{2}$ and ${\rm Var}$
With compromise 1 and 2, we can simplify $E[g(V_{i}A_{s},B_{i})]$ and ${\rm Var}(g(V_{i}A_{s},B_{i}))$ into the following closed-form expressions:
$E[g(V_{i}A_{s},B_{i})]$ $=E[a_{B_{i}}V_{i}A_{s}+b_{B_{i}}]=a_{B_{i}}\Omega_{i}A_{s}+b_{B_{i}}$
${\rm Var}(g(V_{i}A_{s},B_{i}))$ $={\rm Var}(a_{B_{i}}V_{i}A_{s})=a^{2}_{B_{i}}A_{s}^{\top}{\rm Var}_{\Omega_{i}}A_{s}$
$\text{where}\quad{\rm Var}_{\Omega_{i}}={\rm Var}(V_{i})$ =$=\begin{cases}{\rm Var}(V_{ijj})=B_{i}\Omega_{ij}(1-\Omega_{ij})\quad\text{for $j\in\{1,2,\dots,N\}$.}\\ {\rm Var}(V_{ijk})=-B_{i}\Omega_{ij}\Omega_{ik}\quad\text{when $j\neq k$.}\end{cases}$
We estimate the ${\rm Bias}^{2}$ and ${\rm Var}$ in the loss function for one classroom as follows:
${{\rm Bias}}^{2}=$ $\frac{1}{NQ}\sum_{i=1}^{N}(a_{B_{i}}\Omega_{i}A_{s}+b_{B_{i}}-A_{fi})^{\top}(a_{B_{i}}\Omega_{i}A_{s}+b_{B_{i}}-A_{fi})$
${{\rm Var}}=$ $\frac{1}{NQ}tr(\sum_{i=1}^{N}a_{B_{i}}^{2}A_{s}^{\top}{\rm Var}_{\Omega_{i}}A_{s})$
where $N$ is the number of students in the classroom and $Q$ is the number of questions that are asked in the survey ($Q=10$). For one student, the computational complexity for MSE is the same as that for ${\rm Var}_{\Omega_{i}}$ which is $O(N^{2})$. We do it for $N$ students, thus, the total complexity is $O(N^{3})$.
D.2 Downweighing variance
The benefit of downweighing variance can be intuitively explained with social network density. The classroom network exhibits a dense nature: students report more than 1 friend (most of the students report 5 friends which is the maximum number of friends that they can report). When $(\mu,\kappa,\lambda)=(1,0,0)$, we observe that fitted $\Omega$ matrix predicts that all friendship forming probability is concentrated on one boy and one girl, all boys choose one boy as their only friend; all girls choose one girl as their only friend. This happens because extremely concentrated friendship forming probability $\Omega$ results in close to zero variance, indicating that variance is overweighted. Consequently, the neural network predicts an almost deterministic friendship formation rule, with each row having one entry being close to one and other entries being close to zero. This violates the dense nature of classroom social networks, therefore, we downweigh variance to account for social network density.
D.3 Using fitted $\Omega$ to select $(\mu,\kappa,\lambda)$
By leveraging common knowledge of social network structure, we can avoid cross-validation when tuning hyperparameters $(\mu,\kappa,\lambda)$. We examine whether the $\Omega$ matrix generated by a combination of $(\mu,\kappa,\lambda)$ satisfies the commonly known properties of social networks. Such examination uses information from train data exclusively and does not require any validation data.
For example, as mentioned in Appendix D.2, $(\mu,\kappa,\lambda)=(1,0,0)$ is unreasonable as the $\Omega$ matrix generated by this combination of tuning parameters violates the dense nature of classroom network.
Another example is when $(\mu,\kappa,\lambda)=(1,60,100)$, $\Omega$ is almost a constant within each column: all boys are equally influenced by their boy classmates; all girls are equally influenced by their girl classmates. This violates the clustering property of social networks: students should exhibit varying degrees of influence on different groups of students.
We pick $(\mu,\kappa,\lambda)=(0.2,0.3,0.3)$. The choice of hyperparameters might seem ad hoc; however, one can formalize measures of gender homophily, node centrality, and clustering to optimize these hyperparameter choices more systematically. Such a tuning method is time-efficient because we only need to fit the model once for a particular choice of $(\mu,\kappa,\lambda)$ in comparison to $k$-fold cross-validation which requires $k$ instances of the fitting.
Appendix E Linear-in-means model
The conventional econometric model for peer effect measurement is the linear-in-means model
$y_{ics}=\beta\underbrace{\frac{\sum_{j=1,j\neq i}^{N_{cs}}z_{jcs}}{N_{cs}-1}}_{\text{Exclusive means}}+X_{ics}\gamma+\underbrace{\theta_{s}}_{FE}+\underbrace{\mu_{cs}}_{RE}+\epsilon_{ics}$
where $y_{ics}$ is student $i$’s 8th-grade cognitive ability in class $c$ school $s$; $N_{cs}$ is the class size of classroom $c$ from school $s$; $z_{jcs}$ is classmate’s 6th-grade class rank in class $c$ school $s$; $X_{ics}$ is additional controls for student $i$ such as gender, age, ethnic group, and parents’ education; $\theta_{s}$ is school fixed effects and $\mu_{cs}$ is class random effects. We can rewrite linear-in-means model as the $y_{cs}=\beta W_{cs}z_{cs}+X_{cs}\gamma+\theta_{s}+\mu_{cs}+\epsilon_{ics}$ where
$W_{cs}=\begin{pmatrix}0&\frac{1}{N_{cs}-1}&\frac{1}{N_{cs}-1}&\dots&\dots&\frac{1}{N_{cs}-1}\\ \frac{1}{N_{cs}-1}&0&\frac{1}{N_{cs}-1}&\dots&\dots&\frac{1}{N_{cs}-1}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \frac{1}{N_{cs}-1}&\frac{1}{N_{cs}-1}&\frac{1}{N_{cs}-1}&\dots&\dots&0\end{pmatrix}.$
Linear-in-means model implicitly assumes that all students exert the same amount of peer influence on all their classmates as evidenced by the $W_{cs}$ matrix. The structural equation that contains the parameter of interest $\beta$ (i.e. Equation (3)) replaces $W_{cs}$ with $\Omega_{cs}$. The microfoundation for specification of Equation (3) is the following.
E.1 Microfoundation of Eq(3)
In our model, we assume that students continuously update their best friend, and momentarily, a student receives peer effect exclusively from his best friend. Therefore, the peer effect a student receives from one of his classmates is proportional to the probability of him considering that classmate to be his best friend.
As illustrated by example C.4, ${\rm Cov}(\Omega_{cs},\epsilon_{cs})\neq 0$. In economics, this problem is called endogeneity. We address endogeneity by constructing an instrumental variable which happens to be $W_{cs}z_{cs}$.
E.2 Comparison between linear-in-means and friendship-weighted specifications
The two $\beta$s in Eq(2) and Eq(3) have different interpretations. For linear-in-means model, $\beta$ in Eq(2) is the marginal effect of middle school classmates’ average 6th-grade class quantile on the outcome; for friendship-weighted specification, $\beta$ in Eq(3) is the marginal effect of middle school friends’ weighted average 6th-grade class quantile on the outcome. The former has very little policy implication on the average peer effect since the linear-in-means model essentially restricts how much total peer effect each student can exert on other classmates. Our friendship-weighted specification relaxes this restriction: a popular student is allowed to exert more influence on all his classmates than another less popular student. Students may also have different levels of influence on different subgroups, for example, in our empirical results, boys influence boys more and girls influence girls more. This allows us to devise average-maximizing class assignments.
Appendix F Genetic algorithm used to optimize class assignment policy
We constrain that for a school which has fewer girls than boys, it has to assign at least 35% and at most 65% of the girls to one classroom; for a school which has fewer boys than girls, it has to assign at least 35% and at most 65% of the boys to one classroom.
Some of the parameters that we use for our GA design include (1) the number of iterations is 150, (2) the mutation probability is 5%, and (3) the number of swaps for each iteration is 100.
Let $C$ denote the set of all students that the principal is going to divide into classrooms. Let $C_{1}$ and $C_{2}$ denote the set of students from the first and second classrooms, respectively. Note that $(C_{1},C_{2})$ fully pins down the class assignment policy. We constrain that $-1\leq\lvert C_{1}\rvert-\lvert C_{2}\rvert\leq 1$, ensuring that the classrooms are of comparable sizes. For simplicity, we first illustrate two functions that we repeatedly use in our GA design with Algorithm 2 and Algorithm 3.
Algorithm 2 check function
Input: $C_{1},C_{2},N_{b},N_{g},j\in\{b,g\}$
Purpose: Validate whether class assignment meets gender ratio constraints
procedure check($C_{1},C_{2}$)
     if $0.35\cdot N_{j}\leq\sum_{s\in(C_{1})}\mathbbm{1}_{\{s\text{ is gender $j$}\}}\leq 0.65\cdot N_{j}$ then
         return 1
     else
         return 0
     end if
end procedure
Algorithm 3 swap function
Input: $C_{1},C_{2},CP$
Purpose: Swap students between two classrooms
procedure swap($C_{1},C_{2},CP$)
     if $CP\neq({\rm NULL},{\rm NULL})$ then
         $CP_{1},CP_{2}\leftarrow CP$
         $C_{1}\leftarrow(C_{1}\setminus\{CP_{1}\})\cup\{CP_{2}\}$
         $C_{2}\leftarrow(C_{2}\setminus\{CP_{2}\})\cup\{CP_{1}\}$
     end if
     return $C_{1},C_{2}$
end procedure
Appendix G Friendship formation pattern
We show that the friendship formation patterns that we describe in our main text are the same throughout different classroom networks with Figure LABEL:heat_maps_FIM_69_and_138. First, the gender homophily effect is the dominant predictor for friendship formation. Boys almost exclusively make friends with boys and girls make friends with girls. Second, the centrality pattern is clear in the heat maps. Some individual students are extremely influential to other students. Third, students have varying degrees of influence on other students in the same classroom. Even within the same gender, one student is considered by other students as best friends with different probabilities.
In Algorithm 4, we describe how we compute the prediction error for the PeerNN model with test data. In the algorithm, we denote function $G$ as a modification of correspondence $\mathcal{G}$, where if $\mathcal{G}$ outputs multiple outcomes, function $G$ takes the average of them; otherwise, $\mathcal{G}$ and $G$ share the same output. For example, when $V_{i}\cdot A_{s}=1$ and $B_{i}=1$, $\mathcal{G}$ outputs $\{2,3\}$ as shown by Table 5, then $G$ outputs 1.5.
Algorithm 4 Computing prediction error for question $q$
Input: $\Omega,B_{i},A_{sq},A_{fq}$
Output: Prediction error
${\rm PE}_{q}=0$
$R=1000$
for $i$ in $1:N$ do
     $\Xi_{i}\leftarrow\Omega_{i}$
     $V\leftarrow{\rm rep}(0,N)$
     $d\leftarrow{\rm rep}(0,R)$
     for $j$ in $1:R$ do
         for $j$ in $1:B_{i}$ do
              $v\leftarrow\nu\sim{\rm Multinomial}(n=1,{\rm prob}=\Xi_{i})$
              $\Xi_{i}\leftarrow\frac{\Xi_{i}-v\circ\Xi_{i}}{1-v\cdot\Xi_{i}}$
              $V\leftarrow V+v$
         end for
         $d[j]\leftarrow G(V\cdot A_{sq},B_{i})-A_{fq}$
     end for
     ${\rm PE}_{q}\leftarrow{\rm PE}_{q}+d\cdot d$
end for
Algorithm 4 describes how prediction error is computed for a classroom with $N$ students. We sum prediction error over all classrooms in the test set.
Appendix H Class assignment results
We show results for multiple schools. From all figures, GA always severely penalizes a few students in pursuit of the highest average peer effect. In contrast, AFGA generates much more fair class assignment policies. In some cases, readers might be able to visually discern that the average peer effect of AFGA is higher than raw assignment as AFGA tends to have lighter color plots.
Figure 9: Classroom 5 and 6, $\phi=\rho=1$
Figure 10: Classroom 25 and 26, $\phi=\rho=2$
Figure 11: Classroom 45 and 46, $\phi=\rho=1$
Figure 12: Classroom 69 and 70, $\phi=\rho=0.5$
Figure 13: Classroom 73 and 74, $\phi=\rho=0.5$
Figure 14: Classroom 77 and 78, $\phi=\rho=1.5$